10.On May 1, 2018, Meta Computer, Inc., enters into a contract to sell 4,500 units of Comfort Office Keyboard to one of its clients, Bionics, Inc., at a fixed price of $76,500, to be settled by a cash payment on May 1. Delivery is scheduled for June 1, 2018. As part of the contract, the seller offers a 25% discount coupon to Bionics for any purchases in the next six months. The seller will continue to offer a 5% discount on all sales during the same time period, which will be available to all customers. Based on experience, Meta Computer estimates a 50% probability that Bionics will redeem the 25% discount voucher, and that the coupon will be applied to $45,000 of purchases. The stand-alone selling price for the Comfort Office Keyboard is $19.00 per unit.

Required:
1. How many performance obligations are in this contract? (number of performance obligations ___________)
2. Prepare the journal entry that Meta would record on May 1, 2018 using the given information. Then, prepare the journal entry assuming the same facts and circumstances as above, except that Meta gives a 5% discount option to Bionics instead of 25%.

(2.1Record the entry for Meta on May 1, 2018.

2.2Record the entry for Meta on May 1, 2018 assuming that Meta gives a 5% discount option to Bionics instead of 25%)
Answer #1

1. How many performance obligations are in this contract? (number of performance obligations :-

In this case there are two performance obligations:

(i) One is the actual keyboard

(ii) Other is customer option for future discount

2.1Record the entry for Meta on May 1, 2018 :-

Date

General Journal

Debit

Credit

May 1, 2018

Cash

$76500

Deferred Revenue – Keyboards

$72675

Deferred Revenue – discount coupon

$3825

Keyboards: $19 * 4500 = $85500

Option: $45000 * (.25 - .05) * .50 = $4500

Allocation:

Keyboards: [$85500 / ($85500 + $4500) = .95 * $76500 = $72675]

Option: [$4500 / ($85500 + $4500) = .05 * $76500 = $3825]

2.2Record the entry for Meta on May 1, 2018 assuming that Meta gives a 5% discount option to Bionics instead of 25% :-

Date

General Journal

Debit

Credit

May 1, 2018

Cash

$76500

Deferred Revenue – Keyboards

$76500